A rock falls from a tower that is 304 feet high. As it is falling, its height is given by the formula h=304-16t^2. How many seconds(in tenths) will it take for the rock to hit the ground (h=0)? (I believe you just set H=0 and then solve? Is that correct?)
Found 3 solutions by stanbon, mananth, Alan3354:
Answer by stanbon(75887)

Yes, the height will be zero when the rock hits the ground.
-------------
Cheers,
Stan H.

Answer by mananth(16946)

A rock falls from a tower that is 304 feet high. As it is falling, its height is given by the formula h=304-16t^2. How many seconds(in tenths) will it take for the rock to hit the ground (h=0)? (I believe you just set H=0 and then solve? Is that correct?)
Yes
16t^2+h-304 =0
use the formula x1, x2 = -b (+, -) sqrt(b^2-4ac) / 2a
a=16, b=1 c=-304
b^2-4ac= 1-19456
= - 1 + sqrt( 19455) / 2*16
= -1+ 139.48 /32
= 138.48 /32
= 4.32
Similarly find x2 and it will be -4.39
So t=4.32 seconds

Answer by Alan3354(69443)

h=304-16t^2 = 0
16t^2 = 304
t^2 = 19
t =~4.36 seconds